# Leveraging Audio and Text Modalities in Mental Health: A Study of LLMs Performance

### v1

@@ 5.9 Limitations
Comparative Evaluation: Additionally, when comparing our results with existing studies, to our knowledge we have not found any work concerning PTSD severity and multi-class classification tasks on the E-DAIC dataset, which are largely unexplored. For depression severity classification, existing studies often employ different labeling systems, such as using all numbers within the range (e.g., 0-24) as labels, which differs from our methodology. This diversity in approaches makes direct comparisons challenging.
@@
    The E-DAIC dataset used for fine-tuning is both small and imbalanced, with: 86 depression samples, 189 Non-depression samples. This imbalance makes it challenging for the model to learn effectively. With limited "depression" examples, the model struggles to generalize, ultimately failing to surpass the performance achieved by zero-shot (ZS) inference on unprocessed data, where no task-specific fine-tuning was conducted.
@@ 6 Conclusion
This study has systematically evaluated the application of Large Language Models (LLMs) in mental health diagnostics, focusing on depression and PTSD using the E-DAIC dataset. We conducted experiments across a range of tasks to evaluate the comparative effectiveness of text and audio modalities and to investigate if a multimodal approach could enhance diagnostic accuracy. Our research introduced custom metrics, the Modal Superiority Score (MSS) and the Disagreement Resolvement Score (DRS), specifically designed to measure how the integration of modalities impacts model performance. The analysis consistently demonstrated that text modality excelled in most tasks, outperforming audio modality. However, the integration of text and audio modalities often led to improved outcomes, suggesting that combining these modalities could leverage the complementary strengths of each. The MSS and DRS metrics provided insights into the extent of these improvements, highlighting scenarios where multimodal approaches showed promise over single modalities.
While these findings underscore the potential of multimodal approaches for enhancing the performance of LLMs, it is important to emphasize that this study is purely computational. The clinical utility of these methods remains untested and must be validated through rigorous controlled clinical studies to determine their effectiveness and reliability in real-world settings. Moreover, these results are derived from a single dataset (E-DAIC), which is relatively small and may not generalize to broader, more diverse populations or clinical environments. Variability in language, speech patterns, and diagnostic criteria across different datasets or real-world conditions could significantly impact model performance.
7 Acknowledgement

### v2

@@ 5.9 Limitations
Comparative Evaluation: Additionally, when comparing our results with existing studies, to our knowledge, we have not found any work concerning PTSD severity and multi-class classification tasks on the E-DAIC dataset, which are largely unexplored. For depression severity classification, existing studies often employ different labeling systems, such as using all numbers within the range (e.g., 0-24) as labels, which differs from our methodology. This diversity in approaches makes direct comparisons challenging.
@@
    The E-DAIC dataset used for fine-tuning is both small and imbalanced, with 86 depression samples and 189 Non-depression samples. This imbalance makes it challenging for the model to learn effectively. With limited "depression" examples, the model struggles to generalize, ultimately failing to surpass the performance achieved by zero-shot (ZS) inference on unprocessed data, where no task-specific fine-tuning was conducted.
@@ 6 Conclusion
This study has systematically evaluated the application of Large Language Models (LLMs) in mental health diagnostics, focusing on depression and PTSD using the E-DAIC dataset. We conducted experiments across a range of tasks to evaluate the comparative effectiveness of text and audio modalities and to investigate whether a multimodal approach could enhance diagnostic accuracy. Our research introduced custom metrics, the Modal Superiority Score (MSS) and the Disagreement Resolution Score (DRS), specifically designed to measure how the integration of modalities impacts model performance. The analysis consistently demonstrated that text modality excelled in most tasks, outperforming audio modality. However, the integration of text and audio modalities often led to improved outcomes, suggesting that combining these modalities could leverage the complementary strengths of each. The MSS and DRS metrics provided insights into the extent of these improvements, highlighting scenarios where multimodal approaches showed promise over single modalities.
While these findings underscore the potential of multimodal approaches for enhancing the performance of LLMs, it is essential to emphasize that this study is purely computational. The clinical utility of these methods remains untested and must be validated through rigorous controlled clinical studies to determine their effectiveness and reliability in real-world settings. Moreover, these results are derived from a single dataset (E-DAIC), which is relatively small and may not generalize to broader, more diverse populations or clinical environments. Variability in language, speech patterns, and diagnostic criteria across different datasets or real-world conditions could significantly impact model performance.
Acknowledgement